Grade 11Physical Chemistry

What mass of MnO2 is required to react with 250 ml of 0.5 M HCl in the reation given below;MnO2 + 4HCl MnCl2 + 2H2O + Cl2 [at. wt. of Mn = 55, O = 16, Cl=35.5, H=1]

siddhant
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
First we need to calcultae the valence factor for Mn in the given reaction.
Therefore,
v.f. for Mn => (4-2)*1 = 2
Hence,
moles of MnO2 x valence factor = 250*0.5*1/1000
moles = > 125/2*1000 => 0.0625
mass of MnO2= > 0.0625*87 = 5.67 gms.